Staff:

LBTO:Alessandro(on site), Juan Carlos(remote)

Arcegri: Enrico(remote)

Summary

Investigating the lrge drift of the pupil in y direction observed during the rotation wobble check.

it has been observed that the camera lens in the Y-Axis moves all the range from negative to positive with an intrussion in the limits about 8 pixels.

Discussing the last plots from October,22,23 and 25 with Enrico and what it could be the reason of this movements; a few actions were taken to understand this behaviour.
  • First inspect the AGW is something moves.
  • Explore the possible other alternatives in the optical path.
The inspection in the AGW showed that the OCAM has a small pivot movement(pivot).

Details.

Telescope at horizon and access to the AGW. Alessandro access to the AGW.

9:00 am Power On WFS

Alignment of the bayside stage to the internal source:
  • Board setup: telescope_internal_source_101421
Alessandro inspect the OCAM and it showed that the OCAM moves. See the attach video: VID-20211026-WA0004.mp4.

The image below shows the area where it is vible the movements.

tempFileForShare_20211026-131242.jpg

Inspecting if there is any screw loose and it seems that some screw is loose inside of the OCAM.

discussing with Enrico the finds and the next steps.

It will be good to get a reference vector of internal pupil respect to the telescope pupil for the position where the pupils will be displayed on the CCD.

11:00 am Alsessandro doing LBC.

2:00 PM back to AGW.

Alignment of the internal pupils.taking a TN.
  • done.
  • Deploy ArgosCalunit.
  • Set flat the Adsec.
  • CCD47 not live
  • change binning.->CCD 47 OK.
  • internal light off.
  • CCD47 Dark.
Tue Oct 26 21:47:51.749 2021 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[1] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Tue Oct 26 21:47:51.764 2021 aos.info - WFS source set to: FLAOWFS
Tue Oct 26 21:47:52.819 2021 aos.info - FLAO: WFS: loading setup for LUCIFER

preset->OK

Acquire.->Error

2021-10-26 21:52:49.471709| -13006 - [AOException] centerStar: FATAL - [Errno 2] No such file or directory: '/home/AOeng/pause_resume_test_20211011/aoroot/calib/wfs/current/ccd47/backgrounds/bin1/20211026_214836.fits'
2021-10-26 21:53:17.068632| Command canceled

Second preset.

t 26 21:53:26.520 2021 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[3] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Tue Oct 26 21:53:26.536 2021 aos.info - WFS source set to: FLAOWFS
Tue Oct 26 21:53:27.217 2021 aos.info - FLAO: WFS: loading setup for LUCIFER
Tue Oct 26 21:53:27.231 2021 aos.info - FLAO: WFS: configuring CCDs
  • center pupils.
  • rerotator off
  • position zero angle.
  • CL y=-100,x=0
  • TN:20211126_
  • WFS setup: 100Hz,bin1,10l/D,EMC=300
  • TN:220445. not all the pupils illuminated.
  • align pupls using the bayside stage.
  • TN : 221007.
  • close tt loop
  • TN:221904
  • take screenshot
Stop.

3:21 Power off

Adsec rest.

disable rr mode

lookup table enable and temp compensaiton.

Handover 3:30 PM

sent email to Enrico with data to look at.

-- JuanCarlosGuerra - 26 Oct 2021
Topic revision: r3 - 26 Oct 2021, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback